OpenStudy (anonymous):

Which expressions could complete this equation so that it has one solution? 4(4x - 3) - 6x = Choose exactly two answers that are correct. A. 12x - 12 - 2x B. 9x - 12 C. 4x + 2(3x - 6) D. 4(2x + 3) - 24

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

I don't know. Right now I think b and c but I don't know

OpenStudy (anonymous):

@IrishBoy123 he needs help id remember this stuff

OpenStudy (anonymous):

Oh ok cool I really need to get this done and I don't know how to do this

OpenStudy (irishboy123):

question doesn't make any sense to me, i'm afraid there are no equals signs, so there are no equations

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

I think I'll just guess then I'm sure I'll pass even if I miss this one

OpenStudy (irishboy123):

hang on!

OpenStudy (anonymous):

OK

OpenStudy (irishboy123):

there is a pattern in here the question is: 4(4x - 3) - 6x which simplifies to 16x - 12 - 6x = 10x - 12 look at A 12x - 12 - 2x = 10x - 12 so the question is meaningless to me but there is a pattern B 9x - 12 C 4x + 2(3x - 6) = 10x - 12 [again] D 4(2x + 3) - 24 = 8x -12 it's either A&C or B&D depending on what the question means (!!) , if that makes sense....

OpenStudy (irishboy123):

hang on -- you've re-posted the question with an = sign that means that if you have 10x-12 on the RHS, th eequation cancels out so there is no equation so it's B&D

OpenStudy (anonymous):

Thank you so very much!!! Could you check a few of my answers for me please?

OpenStudy (irishboy123):

i gotta go right now, but any other time is good

OpenStudy (anonymous):

Ok well thanks for your help anyways

OpenStudy (anonymous):

https://mathway.com/

OpenStudy (anonymous):

use this for math if you need help

OpenStudy (anonymous):

Thanks